# Certifying the Right to be Forgotten: Primal-Dual Optimization for Sample and Label Unlearning in Vertical Federated Learning

### v1

@@ I Introduction
Federated learning (FL) [mcmahan2017communication, li2020federated, liang2024fedrema] enables collaborative model training without direct data sharing among multiple distributed parties. Nevertheless, AI models can remember sensitive data during the machine learning process, raising privacy concerns related to the “right to be forgotten” established by regulations such as GDPR [regulation2018general]. Such regulations mandate not only the deletion of raw data upon user request but also the removal of its influence on trained models, necessitating solutions capable of removing specific data contributions from machine learning models while maintaining performance. In response, federated unlearning is an attractive approach to address privacy concerns in distributed environments by removing the influence of specific data points from trained models without requiring complete retraining [guo2019certified, liu2024guaranteeing, bourtoule2021machine, cao2015towards, jiang2024feduhb].
@@
To effectively implement federated unlearning, the underlying architecture needs to be considered. Federated learning can be classified into two categories based on data distribution patterns: horizontal federated learning (HFL) [liu2024privacy, boyd2011distributed, jiang2024efficient] where data is partitioned by samples, with each party holding identical features for different user subsets, and vertical federated learning (VFL) [liu2024vertical, yang2019federated, chen2020vafl, zhang2021survey] where data is partitioned by features, with multiple parties training a model on shared users with complementary feature subsets. While horizontal federated unlearning has received significant research attention, unlearning in VFL has garnered less focus [pan2025feature, gu2024few, han2025vertical] and presents greater challenges due to interparty dependencies, especially for sample and label unlearning. In VFL systems, users may request removal of specific data samples to exercise data sovereignty, requiring sample unlearning. Additionally, regulatory policy changes may mandate withdrawal of entire data categories or classes, necessitating label unlearning to remove complete classes along with all associated samples. Despite its importance, this problem remains largely unexplored. Removing samples or labels in VFL affects all participating parties simultaneously, as each party holds different feature subsets of the same samples. This process demands substantial computational and communication resources, while the interdependencies between distributed features complicate the isolation and removal of specific data contributions. Studies such as [varshney2025unlearning, gu2024few] have attempted to address this issue by adopting gradient ascent. However, existing methods often struggle to balance unlearning effectiveness and model performance preservation. Such approaches exhibit instability, sometimes failing to forget target data or causing excessive forgetting that degrades model utility [goodfellow2015empirical, KirkpatrickPRVD16].
@@
To address these challenges, we propose FedORA (Federated Optimization for data Removal via primal-dual Algorithm), adopting a primal-dual optimization framework for sample and label unlearning in VFL. FedORA formulates unlearning as a constrained optimization problem [beck2017first, smith2018cocoa, levy2020large], inspired by Primal-Dual Hybrid Gradient (PDHG). The primal-dual framework naturally balances utility preservation on remaining data with effective forgetting of target data (unlearning data), while leveraging Lagrange duality to mathematically certify unlearning effectiveness. Furthermore, we design an unlearning loss function that encourages uncertainty in classification for unlearning samples rather than pushing the model to misclassify unlearning data, thus avoiding excessive forgetting often associated with gradient ascent methods. In FedORA, stable convergence depends on appropriate step sizes for both primal and dual variable updates. We introduce an adaptive mechanism that tracks how much the model parameters change between successive iterations to dynamically adjust step sizes. Additionally, since the remaining data has already influenced the model during training, processing the entire dataset to maintain model utility during unlearning is unnecessary. FedORA employs an asymmetric batch processing strategy that processes only partial batches of remaining data and full batches of unlearning data, reducing computational overhead. Theoretically, we prove that the difference between the unlearned models obtained by FedORA and Train-from-scratch has an upper bound, providing theoretical validation of FedORA’s unlearning effectiveness. Empirically, we evaluate the approach on the tabular dataset Income and image datasets MedMNIST, CIFAR-10, CIFAR-100, and Tiny-ImageNet across sample and label unlearning scenarios. Results indicate that FedORA achieves unlearning effectiveness comparable to Train-from-scratch with reduced computational and communication costs. Besides, we assess FedORA’s resilience against membership inference attacks and backdoor attacks, validating that our approach effectively removes the influence of unlearning data.
@@ II-A Horizontal Federated Unlearning
In HFL systems, the unlearning target may include either the complete removal of an entire client or specific subsets of data contributed by a client. To achieve unlearning across these varied targets, researchers have developed multiple methods, each offering different advantages depending on the specific implementation context [liu2024survey]. Train-from-scratch (Retrain) rebuilds the model from scratch only using the remaining data after removing the target data [tao2024communication]. While Retrain ensures complete removal of target data contributions, it discards all previous training progress and requires significant computational and communication overhead, making it often impractical for large-scale federated systems. Gradient ascent applies reverse learning by increasing the loss on target data instead of reducing it, effectively reducing the influence of unlearning data [halimi2022federated]. However, gradient ascent can easily lead to forgetting too much information, requiring careful controls to preserve knowledge from remaining data. Fine-tuning calibrates the existing model using only the remaining data to reduce the impact of target data [jiang2024towards]. While more efficient than Retrain, fine-tuning still involves multiple iterations and increased storage costs. Model scrubbing applies mathematical changes to model parameters to approximate a model trained only on remaining data without actual retraining [cao2023fedrecover]. Most scrubbing approaches rely on mathematical approximations, which are computationally challenging for high-dimensional models. Synthetic data methods replace target data with synthetic data to help the model forget specific information [li2023federated, xu2023revocation]. For example, synthetic labels can be generated for target data and used during training to accomplish unlearning.
@@ II-B Vertical Federated Unlearning
Fig. 1 illustrates the fundamental distinctions between three unlearning objectives within vertical federated learning: sample unlearning, label unlearning, and feature unlearning. Sample unlearning removes partial samples from specific classes while preserving other samples. Label unlearning completely removes entire classes along with all associated samples. Feature unlearning addresses the unique VFL scenario by removing feature contributions from specific participating parties regardless of samples or classes. Research on VFU remains in its nascent stages. Deng et al. [deng2023vertical] modify model parameters through inverse gradient operations or parameter adjustments for unlearning. Wang et al. [wang2024efficient] propose to accelerate retraining using optimizer switching (RAdam to SGDM) and checkpoint loading. Pan et al. [pan2025feature] apply feature projection, masking, or transformation techniques to selectively forget feature-based information, while Han et al. [han2025vertical] use adversarial techniques to verify unlearning effectiveness. Varshney et al. [varshney2025unlearning] handle client-level unlearning through contribution isolation, while feature-level unlearning may involve feature masking or projection. Current VFU research primarily focuses on feature unlearning, with limited attention to sample and label unlearning. Gu et al. [gu2024few] and Varshney et al. [varshney2025unlearning] employ gradient ascent to forget specific examples or labels, but suffer from instability issues. Sample and label unlearning remain crucial components within the VFL framework. To address this gap, our research introduces a primal-dual optimization framework that effectively balances unlearning effectiveness, computational efficiency, and model performance preservation.
@@ III-A Workflow of VFL
Prior to training, private set intersection (PSI) [pinkas2018scalable], a cryptographic protocol, is employed to securely identify the common user set among all participating parties while ensuring that private information about non-overlapping users remains protected. This yields the intersection set $\mathcal{U}=\bigcap_{i=1}^{N}\mathcal{U}_{i}$, which contains only the overlapping user IDs. Each party then extracts the corresponding feature matrix $\mathbf{A}_{i}\in\mathbb{R}^{|\mathcal{U}|\times d_{i}}$ to ensure consistent input for collaborative learning.
@@ III-C General Setting of PDHG algorithm
PDHG [sadiev2022communication, chambolle2016introduction, malitsky2018first], also known as the Chambolle-Pock algorithm, has become a widely used optimization tool in machine learning applications. Initially introduced for solving convex-concave saddle-point problems [chambolle2011first], PDHG has proven highly effective for handling optimization tasks involving non-smooth terms and linear operators. In its most general formulation, PDHG addresses the following optimization problem:
@@ IV-B Primal-dual Framework for Unlearning
with $\iota_{\Omega\geq 0}$ being the indicator function of the non-negative orthant. Here, $K\Theta$ represents the linearized coupling from the unlearning constraint, denoting the per-iteration first-order approximation of $\mathcal{L}_{u}$ at the current iterate, i.e., $\langle K\Theta,\Omega\rangle\coloneqq-\Omega\langle\nabla\mathcal{L}_{u}(\Theta^{k}),\Theta-\Theta^{k}\rangle$ [Condatpdhglinear], while we keep the compact notation $K\Theta$ for readability. Meanwhile, $G(\Theta)$ captures the regularized loss on the remaining data, and $F^{*}(\Omega)$ encodes the dual constraint along with the non-negativity indicator. The algorithm then proceeds iteratively with the following updates:
@@
where $\circ$ denotes element-wise multiplication. The step sizes $\sigma$ and $\tau$ satisfy $\sigma\tau L^{2}<1$, with $L$ denoting any global Lipschitz constant (or computable upper bound) of $\nabla\mathcal{L}_{u}(\Theta)$ [Condatpdhglinear].
@@ IV-D Theoretic Analysis
The mini-batch gradient estimation error in FedORA is bounded [bottou2018optimization]:
@@
The gradient of the unlearning loss function $\mathcal{L}_{u}(\Theta)$ is $G$-bounded for all parameters $\Theta$ in the feasible region [beck2017first]:
@@
The dual variable sequence $\{\Omega^{k}\}$ generated by the primal-dual algorithm is uniformly bounded: $\|\Omega^{k}\|_{2}\leq\Omega_{max}$ for the constant $\Omega_{max}>0$ and all $k\geq 0$ [chambolle2011first].
@@ V-A2 Datasets and Models
We evaluate our methods on five datasets with tailored VFL architectures. Income [kohavi1996scaling] is a tabular dataset of 48,842 samples for binary income classification. We use VFL-based MLPs. MedMNIST [yang2023medmnist] contains biomedical images across 9 tissue types. We use PathMNIST, a pathology dataset with 107,180 histopathology images of $28\times 28$ pixels with 9 classes. We adopt a VFL-based ResNet18 [he2016deep] architecture. CIFAR-10 [krizhevsky2009learning] has 60,000 RGB images of $32\times 32$ pixels across 10 classes. We adopt a VFL-based ResNet18 [he2016deep] architecture. CIFAR-100 [krizhevsky2009learning] extends CIFAR-10 with 100 classes. We use a VFL-based DenseNet [huang2017densely] architecture. Tiny-ImageNet [le2015tiny] contains 200 classes with 100,000 training images at $64\times 64$ pixels. We use a VFL-based ResNet50 architecture.
@@ V-A4 Baseline
We compare our proposed approach with four unlearning methods. Train-from-Scratch (Retrain) serves as the standard for unlearning, where the model is completely retrained using only the remaining data after removing the target samples to be forgotten. Gradient ascent (GA) [varshney2025unlearning, gu2024few] realizes sample unlearning in VFL by maximizing the loss on unlearning data. ICO [li2025inverse] computes the contribution of data to be forgotten to model parameters, then subtracts these contributions from current parameters and performs gradient correction optimization. CVFU [wang2025forgetting] achieves asynchronous data unlearning by maintaining a confidence matrix at the active party and updating the confidence values within it.
@@ V-A5 Metrics
Utility Preservation. We measure how well the model maintains its performance on classification tasks after unlearning using accuracy on the test data. Higher test accuracy (Acc.) indicates better utility preservation.
@@
Attack Resilience. To assess model robustness after unlearning, we employ membership inference attacks [shokri2017membership] and backdoor attacks [liu2020reflection] to evaluate whether unlearning samples have been effectively removed.
@@ V-B2 Utility Preservation
Based on the test accuracy comparison in Table IV and Table V, we observe distinct utility preservation patterns across sample and label unlearning scenarios. In sample unlearning scenarios, on the tabular Income dataset, FedORA maintains performance closely matching Retrain, while GA, ICO, and CVFU show moderate accuracy decline compared to Retrain, achieving only around 80%. Moreover, unlearning 25% samples from two classes achieves slightly higher test accuracy than unlearning 50% samples from a single class. For image datasets, performance patterns become more evident with increasing complexity. On MedMNIST, FedORA again approaches Retrain’s performance, while other methods show some decline. As datasets become larger and more complex, all five unlearning methods experience accuracy degradation. On MedMNIST and CIFAR-10, both Retrain and FedORA maintain accuracy above 80%, while on CIFAR-100 and Tiny-ImageNet, performance drops below 70%. Moreover, the performance gap between FedORA and other methods becomes more apparent. GA, ICO, and CVFU exhibit substantial accuracy drops, while FedORA maintains stability close to Retrain. Additionally, increasing the number of unlearning classes worsens performance, with the accuracy gap between different unlearning settings becoming more substantial on complex datasets. Label unlearning presents greater challenges than sample unlearning across all methods and datasets. When unlearning the same number of classes but with 100% sample removal, all methods experience more severe accuracy decline compared to unlearning 50% of samples. When unlearning more classes, accuracy also shows further deterioration. However, FedORA maintains its advantage, closely matching Retrain’s performance and even slightly surpassing Retrain on CIFAR-100 and Tiny-ImageNet.
@@ VI Limitations and Future Work
Model Scale and Scalability. While our work demonstrates effective unlearning in VFL using conventional neural network architectures, extension to large language models (LLM) remains unexplored due to computational and architectural constraints. Recent advances [lin2024splitlora, lin2025hsplitlora, han2024parameter] demonstrate the potential for applying federated learning paradigms to large language models through parameter-efficient fine-tuning techniques. However, adapting our primal-dual unlearning framework to LLM would require addressing additional challenges including parameter-efficient fine-tuning integration and specialized handling of transformer architectures in VFL settings. SplitLoRA [lin2024splitlora] and HSplitLoRA [lin2025hsplitlora] provide valuable insights into how model splitting and LoRA can be combined for efficient LLM training, which could potentially be extended to support unlearning scenarios. In the future work, we can integrate our primal-dual framework with parameter-efficient techniques like LoRA for large-scale model unlearning.

### v2

@@ I Introduction
Federated learning (FL) [40, 29, 31] enables collaborative model training without direct data sharing among multiple distributed parties. Nevertheless, AI models can remember sensitive data during the machine learning process, raising privacy concerns related to the “right to be forgotten” established by regulations such as GDPR [43]. Such regulations mandate not only the deletion of raw data upon user request but also the removal of its influence on trained models, necessitating solutions capable of removing specific data contributions from machine learning models while maintaining performance. In response, federated unlearning is an attractive approach to address privacy concerns in distributed environments by removing the influence of specific data points from trained models without requiring complete retraining [14, 36, 3, 6, 21].
@@
To effectively implement federated unlearning, the underlying architecture needs to be considered. Federated learning can be classified into two categories based on data distribution patterns: horizontal federated learning (HFL) [38, 4, 22] where data is partitioned by samples, with each party holding identical features for different user subsets, and vertical federated learning (VFL) [34, 53, 9, 54] where data is partitioned by features, with multiple parties training a model on shared users with complementary feature subsets. While horizontal federated unlearning has received significant research attention, unlearning in VFL has garnered less focus [41, 13, 16] and presents greater challenges due to interparty dependencies, especially for sample and label unlearning. In VFL systems, users may request removal of specific data samples to exercise data sovereignty, requiring sample unlearning. Additionally, regulatory policy changes may mandate withdrawal of entire data categories or classes, necessitating label unlearning to remove complete classes along with all associated samples. Despite its importance, this problem remains largely unexplored. Removing samples or labels in VFL affects all participating parties simultaneously, as each party holds different feature subsets of the same samples. This process demands substantial computational and communication resources, while the interdependencies between distributed features complicate the isolation and removal of specific data contributions. Studies such as [48, 13] have attempted to address this issue by adopting gradient ascent. However, existing methods often struggle to balance unlearning effectiveness and model performance preservation. Such approaches exhibit instability, sometimes failing to forget target data or causing excessive forgetting that degrades model utility [12, 23].
@@
To address these challenges, we propose FedORA (Federated Optimization for data Removal via primal-dual Algorithm), adopting a primal-dual optimization framework for sample and label unlearning in VFL. FedORA formulates unlearning as a constrained optimization problem [1, 46, 27], inspired by Primal-Dual Hybrid Gradient (PDHG). The primal-dual framework naturally balances utility preservation on remaining data with effective forgetting of target data (unlearning data), while leveraging Lagrange duality to mathematically certify unlearning effectiveness. Furthermore, we design an unlearning loss function that encourages uncertainty in classification for unlearning samples rather than pushing the model to misclassify unlearning data, thus avoiding excessive forgetting often associated with gradient ascent methods. In FedORA, stable convergence depends on appropriate step sizes for both primal and dual variable updates. We introduce an adaptive mechanism that tracks how much the model parameters change between successive iterations to dynamically adjust step sizes. Additionally, since the remaining data has already influenced the model during training, processing the entire dataset to maintain model utility during unlearning is unnecessary. FedORA employs an asymmetric batch processing strategy that processes only partial batches of remaining data and full batches of unlearning data, reducing computational overhead. Theoretically, we prove that the difference between the unlearned models obtained by FedORA and Train-from-scratch has an upper bound, providing theoretical validation of FedORA’s unlearning effectiveness. Empirically, we evaluate the approach on the tabular dataset Income and image datasets MedMNIST, CIFAR-10, CIFAR-100, and Tiny-ImageNet across sample and label unlearning scenarios. Results indicate that FedORA achieves unlearning effectiveness comparable to Train-from-scratch with reduced computational and communication costs. Besides, we assess FedORA’s resilience against membership inference attacks and backdoor attacks, validating that our approach effectively removes the influence of unlearning data.
@@ II-A Horizontal Federated Unlearning
In HFL systems, the unlearning target may include either the complete removal of an entire client or specific subsets of data contributed by a client. To achieve unlearning across these varied targets, researchers have developed multiple methods, each offering different advantages depending on the specific implementation context [37]. Train-from-scratch (Retrain) rebuilds the model from scratch only using the remaining data after removing the target data [47]. While Retrain ensures complete removal of target data contributions, it discards all previous training progress and requires significant computational and communication overhead, making it often impractical for large-scale federated systems. Gradient ascent applies reverse learning by increasing the loss on target data instead of reducing it, effectively reducing the influence of unlearning data [15]. However, gradient ascent can easily lead to forgetting too much information, requiring careful controls to preserve knowledge from remaining data. Fine-tuning calibrates the existing model using only the remaining data to reduce the impact of target data [20]. While more efficient than Retrain, fine-tuning still involves multiple iterations and increased storage costs. Model scrubbing applies mathematical changes to model parameters to approximate a model trained only on remaining data without actual retraining [5]. Most scrubbing approaches rely on mathematical approximations, which are computationally challenging for high-dimensional models. Synthetic data methods replace target data with synthetic data to help the model forget specific information [30, 51]. For example, synthetic labels can be generated for target data and used during training to accomplish unlearning.
@@ II-B Vertical Federated Unlearning
Fig. 1 illustrates the fundamental distinctions between three unlearning objectives within vertical federated learning: sample unlearning, label unlearning, and feature unlearning. Sample unlearning removes partial samples from specific classes while preserving other samples. Label unlearning completely removes entire classes along with all associated samples. Feature unlearning addresses the unique VFL scenario by removing feature contributions from specific participating parties regardless of samples or classes. Research on VFU remains in its nascent stages. Deng et al. [11] modify model parameters through inverse gradient operations or parameter adjustments for unlearning. Wang et al. [50] propose to accelerate retraining using optimizer switching (RAdam to SGDM) and checkpoint loading. Pan et al. [41] apply feature projection, masking, or transformation techniques to selectively forget feature-based information, while Han et al. [16] use adversarial techniques to verify unlearning effectiveness. Varshney et al. [48] handle client-level unlearning through contribution isolation, while feature-level unlearning may involve feature masking or projection. Current VFU research primarily focuses on feature unlearning, with limited attention to sample and label unlearning. Gu et al. [13] and Varshney et al. [48] employ gradient ascent to forget specific examples or labels, but suffer from instability issues. Sample and label unlearning remain crucial components within the VFL framework. To address this gap, our research introduces a primal-dual optimization framework that effectively balances unlearning effectiveness, computational efficiency, and model performance preservation.
@@ III-A Workflow of VFL
Prior to training, private set intersection (PSI) [42], a cryptographic protocol, is employed to securely identify the common user set among all participating parties while ensuring that private information about non-overlapping users remains protected. This yields the intersection set $\mathcal{U}=\bigcap_{i=1}^{N}\mathcal{U}_{i}$, which contains only the overlapping user IDs. Each party then extracts the corresponding feature matrix $\mathbf{A}_{i}\in\mathbb{R}^{|\mathcal{U}|\times d_{i}}$ to ensure consistent input for collaborative learning.
@@ III-C General Setting of PDHG algorithm
PDHG [44, 8, 39], also known as the Chambolle-Pock algorithm, has become a widely used optimization tool in machine learning applications. Initially introduced for solving convex-concave saddle-point problems [7], PDHG has proven highly effective for handling optimization tasks involving non-smooth terms and linear operators. In its most general formulation, PDHG addresses the following optimization problem:
@@ IV-B Primal-dual Framework for Unlearning
with $\iota_{\Omega\geq 0}$ being the indicator function of the non-negative orthant. Here, $K\Theta$ represents the linearized coupling from the unlearning constraint, denoting the per-iteration first-order approximation of $\mathcal{L}_{u}$ at the current iterate, i.e., $\langle K\Theta,\Omega\rangle\coloneqq-\Omega\langle\nabla\mathcal{L}_{u}(\Theta^{k}),\Theta-\Theta^{k}\rangle$ [10], while we keep the compact notation $K\Theta$ for readability. Meanwhile, $G(\Theta)$ captures the regularized loss on the remaining data, and $F^{*}(\Omega)$ encodes the dual constraint along with the non-negativity indicator. The algorithm then proceeds iteratively with the following updates:
@@
where $\circ$ denotes element-wise multiplication. The step sizes $\sigma$ and $\tau$ satisfy $\sigma\tau L^{2}<1$, with $L$ denoting any global Lipschitz constant (or computable upper bound) of $\nabla\mathcal{L}_{u}(\Theta)$ [10].
@@ IV-D Theoretic Analysis
The mini-batch gradient estimation error in FedORA is bounded [2]:
@@
The gradient of the unlearning loss function $\mathcal{L}_{u}(\Theta)$ is $G$-bounded for all parameters $\Theta$ in the feasible region [1]:
@@
The dual variable sequence $\{\Omega^{k}\}$ generated by the primal-dual algorithm is uniformly bounded: $\|\Omega^{k}\|_{2}\leq\Omega_{max}$ for the constant $\Omega_{max}>0$ and all $k\geq 0$ [7].
@@ V-A2 Datasets and Models
We evaluate our methods on five datasets with tailored VFL architectures. Income [24] is a tabular dataset of 48,842 samples for binary income classification. We use VFL-based MLPs. MedMNIST [52] contains biomedical images across 9 tissue types. We use PathMNIST, a pathology dataset with 107,180 histopathology images of $28\times 28$ pixels with 9 classes. We adopt a VFL-based ResNet18 [18] architecture. CIFAR-10 [25] has 60,000 RGB images of $32\times 32$ pixels across 10 classes. We adopt a VFL-based ResNet18 [18] architecture. CIFAR-100 [25] extends CIFAR-10 with 100 classes. We use a VFL-based DenseNet [19] architecture. Tiny-ImageNet [26] contains 200 classes with 100,000 training images at $64\times 64$ pixels. We use a VFL-based ResNet50 architecture.
@@ V-A4 Baseline
We compare our proposed approach with four unlearning methods. Train-from-Scratch (Retrain) serves as the standard for unlearning, where the model is completely retrained using only the remaining data after removing the target samples to be forgotten. Gradient ascent (GA) [48, 13] realizes sample unlearning in VFL by maximizing the loss on unlearning data. ICO [28] computes the contribution of data to be forgotten to model parameters, then subtracts these contributions from current parameters and performs gradient correction optimization. CVFU [49] achieves asynchronous data unlearning by maintaining a confidence matrix at the active party and updating the confidence values within it.
@@ V-A5 Metrics
Utility Preservation. We measure how well the model maintains its performance on classification tasks after unlearning using accuracy on the remaining data. Higher test accuracy (Acc.) indicates better utility preservation.
@@
Attack Resilience. To assess model robustness after unlearning, we employ membership inference attacks [45] and backdoor attacks [35] to evaluate whether unlearning samples have been effectively removed.
@@ V-B2 Utility Preservation
Based on the test accuracy on the remaining data in Table IV and Table V, we observe distinct utility preservation patterns across sample and label unlearning scenarios. In sample unlearning scenarios, on the tabular Income dataset, FedORA maintains performance closely matching Retrain, while GA, ICO, and CVFU show moderate accuracy decline compared to Retrain, achieving only around 80%. Moreover, unlearning 25% samples from two classes achieves slightly higher test accuracy than unlearning 50% samples from a single class. For image datasets, performance patterns become more evident with increasing complexity. On MedMNIST, FedORA again approaches Retrain’s performance, while other methods show some decline. As datasets become larger and more complex, all five unlearning methods experience accuracy degradation. On MedMNIST and CIFAR-10, both Retrain and FedORA maintain accuracy above 80%, while on CIFAR-100 and Tiny-ImageNet, performance drops below 70%. Moreover, the performance gap between FedORA and other methods becomes more apparent. GA, ICO, and CVFU exhibit substantial accuracy drops, while FedORA maintains stability close to Retrain. Additionally, increasing the number of unlearning classes worsens performance, with the accuracy gap between different unlearning settings becoming more substantial on complex datasets. Label unlearning presents greater challenges than sample unlearning across all methods and datasets. When unlearning the same number of classes but with 100% sample removal, all methods experience more severe accuracy decline compared to unlearning 50% of samples. When unlearning more classes, accuracy also shows further deterioration. However, FedORA maintains its advantage, closely matching Retrain’s performance and even slightly surpassing Retrain on CIFAR-100 and Tiny-ImageNet.
@@ VI Limitations and Future Work
Model Scale and Scalability. While our work demonstrates effective unlearning in VFL using conventional neural network architectures, extension to large language models (LLM) remains unexplored due to computational and architectural constraints. Recent advances [32, 33, 17] demonstrate the potential for applying federated learning paradigms to large language models through parameter-efficient fine-tuning techniques. However, adapting our primal-dual unlearning framework to LLM would require addressing additional challenges including parameter-efficient fine-tuning integration and specialized handling of transformer architectures in VFL settings. SplitLoRA [32] and HSplitLoRA [33] provide valuable insights into how model splitting and LoRA can be combined for efficient LLM training, which could potentially be extended to support unlearning scenarios. In the future work, we can integrate our primal-dual framework with parameter-efficient techniques like LoRA for large-scale model unlearning.